OpenStudy (anonymous):

Complete the following tasks for this equation:b/7.8 = -2.15. Give an estimate for the solution. Be sure to explain how you arrived at the estimate, either by showing an equation or in words. Solve the equation. Show or explain your work. Explain if your answer is reasonable compared to your estimate.

OpenStudy (anonymous):

You can estimate by approximating the numbers, the equation is about the same as b/8 = -2, so the answer would be about -16 if you multiply by 8 on both sides.

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

The solution is found by multiplying by 7.8 on both sides: b = -2.15 * 7.8 = -16.77

OpenStudy (anonymous):

So the answer is reasonable because it's "pretty close" to -16.

OpenStudy (anonymous):

THANK YOU SO MUCH :)

OpenStudy (anonymous):

No problem! Would you mind pressing "best response" on one of the responses? Glad I could help!